How to factorize it √3*x^2+11*x+6*√3
Sol: √3*x^2+2*x+9*x+6*√3
But then?

Found 3 solutions by Boreal, josgarithmetic, Edwin McCravy:
Answer by Boreal(15235)

√3*x^2+11*x+6*√3=
If you rewrite as sqrt (3) x^2 +2x + 9x +6 sqrt (3)
=x (x sqrt (3)+2) ;; +;; 3 sqrt(3) (x sqrt (3) +2)
=(x+3 sqrt (3))* {x sqrt (3) +2}
=x^2* sqrt(3) +2x+3x(sqrt(3)(sqrt(3) +6 sqrt (3)

Answer by josgarithmetic(39616)

After trying testing roots as for "Rational Roots Theorem", and also trying general solution for quadratic equation, the quadratic expression seems NOT factorable. Discriminant for that expression is -95, therefore the roots are not Real.

Answer by Edwin McCravy(20054)

Boreal's solution is better than the way I did it, so I deleted mine. :)